# LibMoE: A Library for Comprehensive Research Mixture of Experts in Large Language Models

### v2

@@ C.4 Load Balancing Loss and Router Z-loss Experiment
Appendix D Analysis Detailed
D.1 Expert Entropy Allocation
@@
D.2 Router Saturation
@@
D.3 Expert Change Rate
Expert Change Rate (ECR) measures how stable the expert assignments are during training. Given a dataset $\mathcal{D}$}, we compare the expert choices for each token between two checkpoints at times $T$ and $T^{\prime}$. The ECR is the fraction of tokens whose expert assignments have changed.
A low ECR indicates that the router has stabilized, consistently assigning tokens to the same experts. In contrast, a high ECR suggests the router is still changing its policy—signaling instability, ongoing exploration, or lack of convergence. Monitoring ECR helps assess how reliably the routing mechanism is specializing experts over time.
D.4 Router Margin
$\text{Router Margin}(l)=\frac{1}{N}\sum_{i=1}^{N}top1(x_{i})-top2(x_{i}),$
Where:
  • $\text{RouterMargin}(l)$: Router margin at layer $l$
  • NN: The total number of tokens in the dataset.
  • $x_{i}$: $i$-th input token
  • $top1(x_{i})$: top-1 routing score of input $x_{i}$
  • $top2(x_{i})$: top-2 routing score of input $x_{i}$
D.5 Experts Co-activation
@@
Appendix E Training Time and Resource Allocation
Table 9 summarizes the training time and resource utilization across all experimental settings.
@@ Appendix E Training Time and Resource Allocation
Model
Training Time
Resource
Allocation
VLM
Pre-Training 2h35m 4xH100
@@
Visual Instruction
Tuning
SMoE 17h19m 4xH100
@@
Visual Instruction
Tuning
SMoE 33h01m 4xH100
@@
Language
Modeling
158M parametes Dense 5h47m 4xH100
@@
Appendix F Training Benchmark Curves
Figure 18: Benchmark curves during training in language modeling tasks for models with 158M parameters.
Figure 19: Benchmark curves during training in language modeling tasks for models with 679M parameters.

### v3

@@ D.4 Load Balancing Loss and Router Z-loss Experiment
Appendix E Definitions and Formulations of Analysis Metrics
E.1 Total Score Change under Routing Perturbations: DropTop1 vs. DropTop1&2
Let $\mathcal{D}=\{D_{i}\}_{i=1}^{|\mathcal{D}|}$ denote the set of evaluation benchmarks. For a given MoE model, let $M_{i}\in\mathbb{R}$ be its original evaluation score (e.g., accuracy; higher is better) on benchmark $D_{i}$. After applying a routing perturbation $\pi\in\{\textit{DropTop1},\,\textit{DropTop1$\&$2}\}$ at inference time, we obtain the perturbed score $\hat{M}_{i}^{\,\pi}\in\mathbb{R}$.
We define the aggregate performance change (reported as $\Delta$ Total Score in Figure 4) as:
$\Delta\text{TotalScore}^{\,\pi}=\sum_{i=1}^{|\mathcal{D}|}\left(\hat{M}_{i}^{\,\pi}-M_{i}\right).$ (3)
Negative values indicate that the routing perturbation leads to a net performance degradation relative to the original routing configuration, whereas positive values indicate that the perturbed routing achieves higher aggregate performance than the original model. In the latter case, the improvement suggests that the original routing decision may have been suboptimal, and that alternative expert assignments can better exploit the model’s representational capacity.
E.2 Expert Entropy Allocation
@@
E.3 Router Margin
$\text{Router Margin}(l)=\frac{1}{N}\sum_{i=1}^{N}top1(x_{i})-top2(x_{i}),$
Where:
  • $\text{RouterMargin}(l)$: Router margin at layer $l$
  • NN: The total number of tokens in the dataset.
  • $x_{i}$: $i$-th input token
  • $top1(x_{i})$: top-1 routing score of input $x_{i}$
  • $top2(x_{i})$: top-2 routing score of input $x_{i}$
E.4 Router Saturation
@@
E.5 Experts Co-activation
@@
Appendix F Training Time and Resource Allocation
Table 11 reports the training time and GPU resource consumption across all experimental configurations. In the VLM setting, SharedE-V2 and SharedE-V3 consistently demonstrate superior training efficiency, completing training on the OneVision dataset approximately two hours faster than all competing methods. This speedup stems from the use of shared experts, which do not require per-token routing probability computation. As a result, the router only performs inference over the remaining non-shared experts—i.e., $N$ minus the number of shared experts thereby reducing routing overhead and improving overall training efficiency. Importantly, this reduction in runtime does not come at the expense of model quality. As shown in Table 1, both variants remain highly competitive and, in several cases, achieve the best overall performance on OneVision. For language model pretraining, we adopt the CVMM Triton kernel Csordás et al. (2024), which significantly accelerates sparse MoE computation. Consequently, the pretraining runtime in this setting differs substantially from that of the VLM experiments and is dominated by kernel-level optimizations rather than architectural design choices such as SharedE-V2 or SharedE-V3. For this reason, we recommend interpreting training-time comparisons primarily within the VLM setting, where runtime differences more directly reflect methodological design decisions and thus provide a more informative and meaningful comparison.
@@ Appendix F Training Time and Resource Allocation
Model Training Time (hours) Resource Allocation
VLM Pre-Training 2h35m 4xH100
@@
Visual Instruction Tuning LLAVA-665K SMoE 17h19m 4xH100
@@
Visual Instruction Tuning OneVision / 1M2 samples SMoE 33h01m 4xH100
@@
Language Modeling 0.15B parametes Dense 5h47m 4xH100
@@
Appendix G Training Benchmark Curves
Figure 18: Benchmark curves during training in language modeling tasks for models with 0.15B parameters.
Figure 19: Benchmark curves during training in language modeling tasks for models with 0.68B parameters.